OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition. passes through (5, 14), parallel to the line that passes through (12, 2) and (35, 19)

OpenStudy (anonymous):

1. Find the slope of the line made by points (12,2) and (35,19) using teh slope formula. 2. That slope is also the slope of the line that passes through (5,14) as parallel lines have teh same slope. 3. Plug in the slope into y=mx+b which is the formula for a line. Then plug in (5,14) into y=mx+b. Solve for b. 4. Now plug in the slope and b (The y-intercept) into y=mx+b for the equation for the line that passes through (5,14) and is parallel to (12,2) and (35,19)

OpenStudy (anonymous):

y = 17/23x + 237/23 right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

To check teh answer, just plug int (5,14).

OpenStudy (anonymous):

Alright thanks so much!